OpenStudy (anonymous):

Prove: sec x(1 + cos x) = 1 + sec x

OpenStudy (anonymous):

see secx*cosx=1

OpenStudy (luigi0210):

distribute!

Parth (parthkohli):

\[\sec(x) = \dfrac{1}{\cos(x)}\]So\[\dfrac{1}{\cos(x)}\left(1 + \cos(x)\right) = 1 + \dfrac{1}{\cos(x)}\]Distribute\[\dfrac{1}{\cos(x)} + 1 = 1 + \dfrac{1}{\cos(x)}\]Don't need to go further

Parth (parthkohli):

The left hand side is just the right hand side, right?

OpenStudy (anonymous):

Presumably

Parth (parthkohli):

So we're done.

Parth (parthkohli):

Could be shorter by memorizing \(\sec(x) \times \cos(x) = 1\), but bleh :p

OpenStudy (anonymous):

wouldn't you prove it like: secx+secx*cosx =secx+1/cosx*cosx =secx+1

Parth (parthkohli):

Yeah, that's the shorter way to prove it.

OpenStudy (anonymous):

:D

OpenStudy (jhannybean):

can you elaborate @isuckatmath9999 ? :)i mean elaborate your way. haha

OpenStudy (zzr0ck3r):

lol there are about 5000* ways, anyone else got another?

Parth (parthkohli):

@Jhannybean \(\sec(x) \times \cos(x) = 1\)

OpenStudy (anonymous):

I don't neeed another its for an assignment lol. :)

OpenStudy (jhannybean):

yeahhh but how'd he get...sec (x) + 1/cos^2 (x)...

OpenStudy (jhannybean):

OH YOU FACTORED OUT THE SEC (X)???

OpenStudy (anonymous):

yea lol

Parth (parthkohli):

It's not 1/cos^2(x) lol It's 1/cos(x) * cos(x)

OpenStudy (jhannybean):

oh ho ho, i gotchu.

OpenStudy (anonymous):

I have more questions to debate upon lol. don't worry :P heheee

Parth (parthkohli):

I haven't done trig in months. Wow

OpenStudy (anonymous):

there isn't one

OpenStudy (jhannybean):

i feel so dyslexic...